OpenStudy (anonymous):

Solve. 8 = 2^x + 4 ? x = -4 x = -1 x = 0 x = 7

OpenStudy (anonymous):

\[{2^3} ={2^x+4}\\ {3}={x+4}\\ {3-4 }= { x}\\ {x }= { -1} \]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

kk